OpenStudy (anonymous):

the product of two numbers is 88. algebraically determine the number if onenumber is 3 is more than the other

OpenStudy (matt101):

Say the two numbers are x and y. That means xy=88. If y is three more than x, then y=x+3. That means x(x+3)=88. x(x+3)=88 x^2 + 3x = 88 x^2 + 3x - 88 = 0 (x-8)(x+11)=0 x = 8, -11 Therefore the numbers could be 8 and 11 or -11 and -8. Either way, one is 3 more than the other.